# A proof of Xin-Zhang’s tridiagonal determinant conjecture

### v1

@@ Abstract.
We confirm a recent conjecture by Xin and Zhang, which establishes a simple product formula for the characteristic polynomial of an $(n-1)\times(n-1)$ tridiagonal matrix $C$. This characteristic polynomial arises from a recurrence relation that enumerates $n\times n$ nonnegative integer matrices with all row and column sums equal to $t$, also called the Ehrhart polynomial of the $n$th Birkhoff polytope. The proof relies on an unexpected observation: shifting $C$ by a scalar multiple of the identity matrix yields a matrix similar to a lower triangular matrix. In triangular form, the characteristic polynomial reduces to the product of the diagonal entries, leading to the desired closed-form expression. Moreover, we extend this method to broader families of tridiagonal matrices. This provides a new approach for deriving exact enumeration formulas.
@@
Keywords: determinant; tridiagonal matrix; characteristic polynomial; combinatorial identity.
@@ 1. Introduction
Our main objective in this paper is to prove the following conjecture of Xin and Zhang [13], which gives a product formula for the characteristic polynomial of a certain tridiagonal matrix.
Conjecture 1.1 ([13, Conjecture 4.8]).
@@
This conjecture arises in the study of the Ehrhart polynomial $H_{n}(t)$ of the $n$th Birkhoff polytope, where $H_{n}(t)$ counts the number of $n\times n$ nonnegative integer matrices with all row and column sums equal to $t$. The computation of $H_{n}(t)$ has been a subject of extensive research; see, for example, [1, 2, 3, 5, 6, 7, 8, 9, 10, 11, 14]. So far, $H_{n}(t)$ is known only for $n\leq 9$, with $H_{9}(t)$ computed by Beck and Pixton [4] using residue techniques.
Following [4, Corollary 4], it is known that
@@
The characteristic polynomial of $\widetilde{C}^{(n)}$ is still hard: none of the minors of $tI-\widetilde{C}^{(n)}$ admit a simple factorization. This prevents us from proving the conjecture by traditional techniques such as diagonalization, LU decompositions, continued fractions, etc.. Ultimately, we discovered that $\widetilde{C}^{(n)}$ is similar to a lower triangular matrix. We state this result as follows.
@@
The rest of this paper is structured as follows. In Section 2, we prove Theorem 1.2 using an explicit formula of $U^{-1}$, together with several combinatorial identities. In Section 3, we extend these ideas to prove a natural generalization of Theorem 1.2.
@@
To prove Theorem 1.2, we first give a formula of $U^{-1}$, which should be known but we cannot find a reference.
Lemma 2.1.
For a positive integer $n$, let $U$ be the matrix with entries as in (1.3). Then the entries of $U^{-1}$ are given by
(2.4) $(U^{-1})_{i,j}=(-1)^{i+j}U_{i,j}=(-1)^{i+j}\binom{n-i}{n-j}$.
Proof.
Suppose $V$ is the matrix whose entries are given by (2.4). Consider the $(i,j)$-entry of $UV$:
$(UV)_{i,j}=\sum_{l=1}^{n}U_{i,l}V_{l,j}=\sum_{l=1}^{n}(-1)^{l+j}\binom{n-i}{n-l}\binom{n-l}{n-j}$.
It is clear that the above formula equals 0 when $i>j$. When $i\leq j$, we can further write
$(UV)_{i,j}$ $=\sum_{l=i}^{j}(-1)^{l+j}\binom{n-i}{n-l}\binom{n-l}{n-j}$
@@
$=\binom{n-i}{n-j}(1+(-1))^{j-i}=\begin{cases}1,&\mbox{if }i=j\\ 0,&\mbox{if }i<j.\end{cases}.$
Thus, $UV=I$. This completes the proof. ∎
Now, we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Consider the $(i,j)$-entry of $U\widetilde{C}^{(n)}U^{-1}$:
$(U\widetilde{C}^{(n)}U^{-1})_{i,j}$ $=\sum_{1\leq l,m\leq n}(-1)^{m+j}\binom{n-i}{n-l}\binom{n-m}{n-j}(\widetilde{C}^{(n)})_{l,m}$
$=\sum_{l=1}^{n-1}(-1)^{j+l+1}\binom{n-i}{n-l}\binom{n-l-1}{n-j}(n+1-l)(n-l)$
$\qquad+\sum_{l=1}^{n}(-1)^{j+l}\binom{n-i}{n-l}\binom{n-l}{n-j}(n+1-l)(n+1-2l)$
$\qquad+\sum_{l=2}^{n}(-1)^{j+l-1}\binom{n-i}{n-l}\binom{n-l+1}{n-j}(n+1-l)(1-l)$
$=\sum_{l=1}^{n}(-1)^{j+l}\binom{n-i}{n-l}(n+1-l)$
$\hskip 27.74982pt\cdot\Big((n-l)\Big(\binom{n-l}{n-j}-\binom{n-l-1}{n-j}\Big)-(1-l)\Big(\binom{n-l+1}{n-j}-\binom{n-l}{n-j}\Big)\Big)$
$=\sum_{l=1}^{n}(-1)^{j+l}\binom{n-i}{n-l}(n+1-l)$
@@
$=\sum_{l=1}^{n}(-1)^{j+l}\binom{n-i}{n-l}(n+1-l)$
[…]
$=j\sum_{l=1}^{n}(-1)^{j+l}\binom{n-i}{n-l}\binom{n-l}{n-j-1}(n+1-l).$
It is clear that the above formula equals 0 when $i>j+1$. When $i\leq j+1$, we can further write
$(U\widetilde{C}^{(n)}U^{-1})_{i,j}$ $=j\sum_{l=i}^{j+1}(-1)^{j+l}\binom{n-i}{n-l}\binom{n-l}{n-j-1}(n+1-l)$
[…]
@@
$=\begin{cases}i(n-i),&\mbox{if }i=j;\\ (1-i)(n+1-i),&\mbox{if }i=j+1;\\ 0,&\mbox{if }i<j.\end{cases}$
The proof is completed. ∎
@@
And hence $(UMU^{-1})_{i,i-1}=M_{i,i-1}$. The proof is completed. ∎
@@
Observe that $(\widetilde{C}^{(n)})_{i,i}=(\widetilde{C}^{(n)})_{i,i-1}+(\widetilde{C}^{(n)})_{i,i+1}$ for each $i$. This leads us to consider matrix $A$ satisfying $A_{i,i}=r+A_{i,i-1}+A_{i,i+1}$, where $r$ is a parameter. Further exploration gives Theorem 3.1.
@@
Given a positive integer $n$, denote by
@@
We prove the theorem by induction on $n$. It is trivial for $n=1,2$. Assume the theorem holds for n1n-1. We perform elementary operations on $A$ as follows:
@@
Formula (3.5) holds for $1\leq i\leq n-1$, and it is easy to see that $c_{i}^{(s)}=0$ if s+i>n1s+i>n-1. Performing the above elementary operations on $A$ is equivalent to computing the matrix product $PAQ$. Thus, we have

### v2

@@ Abstract.
We confirm a recent conjecture of Xin and Zhang, which establishes a simple product formula for the characteristic polynomial of an $(n-1)\times(n-1)$ tridiagonal matrix $C$. This characteristic polynomial arises from a recurrence relation that enumerates $n\times n$ nonnegative integer matrices with all row and column sums equal to $t$, also called the Ehrhart polynomial of the $n$th Birkhoff polytope. Moreover, we extend our method to broader families of tridiagonal matrices.
@@
Keywords: determinant; tridiagonal matrix; characteristic polynomial; combinatorial identity; Pascal’s triangle.
@@ 1. Introduction
Our main objective in this paper is to prove the following conjecture of Xin and Zhang [12], which gives a product formula for the characteristic polynomial of certain tridiagonal matrix.
Conjecture 1.1 ([12, Conjecture 4.8]).
@@
This conjecture arises in the study of the Ehrhart polynomial $H_{n}(t)$ of the $n$th Birkhoff polytope, where $H_{n}(t)$ counts the number of $n\times n$ nonnegative integer matrices with all row and column sums equal to $t$. The computation of $H_{n}(t)$ has been a subject of extensive research; see, for example, [1, 5, 6, 7, 8, 9, 10, 13]. So far, $H_{n}(t)$ is known only for $n\leq 9$, with $H_{9}(t)$ computed by Beck and Pixton [3] using residue techniques.
Following [3, Corollary 4], it is known that
@@
It is still hard to determine the characteristic polynomial of $\widetilde{C}^{(n)}$: none of the minors of $tI-\widetilde{C}^{(n)}$ admit a simple factorization. This prevents us from proving the conjecture by traditional techniques such as diagonalization, LU decompositions, continued fractions, etc.. Ultimately, we discovered that $\widetilde{C}^{(n)}$ is similar to a lower triangular matrix. We state this result as follows.
@@
The rest of this paper is Section 2, in which we prove Theorem 1.2 using an explicit formula of $U^{-1}$, together with several combinatorial identities. In Section 3, we extend these ideas to prove a natural generalization of Theorem 1.2.
@@
Proof of Theorem 1.2.
By [4, Theorem 1], the entries of $U^{-1}$ are given by
$(U^{-1})_{i,j}=(-1)^{i+j}U_{i,j}=(-1)^{i+j}\binom{n-i}{n-j}$.
Consider the $(i,j)$-entry of $U\widetilde{C}^{(n)}U^{-1}$:
$(U\widetilde{C}^{(n)}U^{-1})_{i,j}$ $=\sum_{1\leq l,m\leq n}U_{i,l}(\widetilde{C}^{(n)})_{l,m}(U^{-1})_{m,j}$
$=\sum_{l=1}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l}(U^{-1})_{l,j}+\sum_{l=2}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l-1}(U^{-1})_{l-1,j}$
$\quad+\sum_{l=1}^{n-1}U_{i,l}(\widetilde{C}^{(n)})_{l,l+1}(U^{-1})_{l+1,j},$
where the second equality holds by $(\widetilde{C}^{(n)})_{l,m}=0$ when |lm|>1|l-m|>1. Observe that $(\widetilde{C}^{(n)})_{l,l}=(\widetilde{C}^{(n)})_{l,l-1}+(\widetilde{C}^{(n)})_{l,l+1}$, and the formulas of $(\widetilde{C}^{(n)})_{l,l-1}$ and $(\widetilde{C}^{(n)})_{l,l+1}$ vanish if setting $l=1$ and $l=n$, respectively. The computation is naturally divided into two parts, denoted by $S_{1}$ and $S_{2}$, respectively. In detail,
$S_{1}$ $=\sum_{l=1}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l-1}(U^{-1})_{l,j}+\sum_{l=2}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l-1}(U^{-1})_{l-1,j}$
$=\sum_{l=1}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l-1}((U^{-1})_{l,j}+(U^{-1})_{l-1,j})$
$=\sum_{l=1}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l-1}(-1)^{j+l-1}\Big(\binom{n-l+1}{n-j}-\binom{n-l}{n-j}\Big)$
@@
Similarly, we have
$S_{2}$ $=\sum_{l=1}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l+1}(U^{-1})_{l,j}+\sum_{l=1}^{n-1}U_{i,l}(\widetilde{C}^{(n)})_{l,l+1}(U^{-1})_{l+1,j}$
$=\sum_{l=1}^{n}U_{i,l}(\widetilde{C}^{(n)})_{l,l+1}(-1)^{j+l}\binom{n-l-1}{n-j-1}$.
Using Equation (2.2), we can further compute $S_{2}$ as:
$S_{2}$ $=\sum_{l=1}^{n}U_{i,l}(-1)^{j+l}(n+1-l)(n-l)\binom{n-l-1}{n-j-1}$
$=\sum_{l=1}^{n}U_{i,l}(-1)^{j+l}(n+1-l)(j-l+1)\binom{n-l}{n-j-1}$.
Then,
$(U\widetilde{C}^{(n)}U^{-1})_{i,j}=S_{1}+S_{2}$ $=j\sum_{l=1}^{n}(-1)^{j+l}\binom{n-i}{n-l}\binom{n-l}{n-j-1}(n+1-l)$
@@
It is clear that the above formula equals 0 when $i>j+1$ and equals (1i)(n+1i)(1-i)(n+1-i) when $i=j+1$. When $i\leq j$, we can further write
$(U\widetilde{C}^{(n)}U^{-1})_{i,j}$ $=j\sum_{l=i}^{j+1}(-1)^{j+l}\binom{n-i}{n-j-1}\binom{j+1-i}{j+1-l}((n-j)+(j+1-l))$
@@
$=\begin{cases}i(n-i),&\mbox{if }i=j;\\ 0,&\mbox{if }i<j.\end{cases}$
This completes the proof. ∎
@@
And hence $(UMU^{-1})_{i,i-1}=M_{i,i-1}$. This completes the proof. ∎
@@
The observation $(\widetilde{C}^{(n)})_{i,i}=(\widetilde{C}^{(n)})_{i,i-1}+(\widetilde{C}^{(n)})_{i,i+1}$ leads us to consider matrix $A$ satisfying $A_{i,i}=r+A_{i,i-1}+A_{i,i+1}$, where $r$ is a parameter. Further exploration gives Theorem 3.1.
@@
Given a positive integer $n$, let
@@
We prove the theorem by induction on $n$. It is trivial for $n=1$ or $2$. Assume the theorem holds for n1n-1. We perform elementary operations on $A$ as follows:
@@
Formula (3.5) holds for $1\leq i\leq n-1$, and it is easy to see that $c_{i}^{(s)}=0$ if s+i>n1s+i>n-1. We further set $c_{i}^{(s)}=0$ for $i\leq 0$. Performing the above elementary operations on $A$ is equivalent to computing the matrix product $PAQ$. Thus, we have